OpenStudy (anonymous):

WILL MEDAL the equation below represents which type of conic section? (x-2)^2/3+(y+2)/4=1 a) parabola b) ellipse c) hyperbola d) circle

OpenStudy (anonymous):

\[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] is an ellipse

OpenStudy (anonymous):

oh is the y NOT squared?

OpenStudy (anonymous):

is it \[\frac{(x-2)^2}{3}+\frac{(y+2)^2}{4}=1\] or is it \[\frac{(x-2)^2}{3}+\frac{y+2}{4}=1\]

OpenStudy (anonymous):

THE SECOND ONE @satellite73

OpenStudy (anonymous):

oh then it is a trick question this is a parabola disguised as an ellipse if that term is not squared it is a parabola